Information for Math 410—Matrix Analysis 1st Summer Session, 1996

Instructor: Doug Ulmer  Class: M-F 1:00-2:45, Engr 311
Office: Math 716  Text:“Matrix Methods” by R. Bronson
Phone: 621-6861
Office Hours: Mondays and Wednesdays 3-4 in my office, or by appointment

    Website: http://www.math.arizona.edu/people/~ulmer/

                teaching/LASm96/LASummer96.html

The course: This course is an introduction to vectors and matrices meant mainly for engineers. The main points are:

(i) row reduction and its applications (finding the solution of systems of linear equations in several variables; calculating inverses and determinants; determining linear dependence or independence and spans of sets of vectors).

(ii) Understanding and calculating eigenvectors and eigenvalues.

(iii) Inner products, orthogonality, projections, and their applications.

(iv) Calculus of matrices. Requirements: The basic requirements are to attend class, to do the homework assignments, and to take the 3 midterm exams and the final exam. Attendance: Experience shows that very few students can succeed in this course without

coming to class every day.

Homework: There will a homework assignment for each day of class. You are required to do the homework, but you are not required to turn it in. (Answers will be provided so you can check your work.) If you don’t do the homework, you will almost certainly not be able to pass the class.

Exams: There will be 3 “midterm” exams, during the last 50 minutes of class on 6/11, 6/18, 6/25 (Tuesdays). The final exam will occupy all of the last class, Wednesday 7/3. The midterm exams and half of the final will cover the preceding week’s material; the other half of the final is over the whole course. There will be no make-up tests.

Grades: Grades will be based entirely on exam scores, with the midterm exams counting 20% each and the final 40%. Course grades will be determined by a straight percentage scale: 90-100% is an A, 80-90% is a B, 70-80% is a C, and 60-70% is a D.

Technology: There are a number of good software packages for doing linear algebra (i.e., working with matrices), in particular the “Linalg” package. Many scientific calculators also have linear algebra functions. There is an open access computer lab in Math 226/228 (open M-Th 9-7 and F 9-5). You can copy and use any of the programs loaded on those machines (including Linalg) for free. These programs run on all standard DOS machines.

Getting help: Feel free to come to office hours if you have questions. Also, there is a free tutoring service in room 320 of the math department from 12 to 3 M-F.

Working together: Please do. You are free to organize yourselves into groups and discuss homework, study for tests, etc. Just make sure that you understand the mathematics for yourself, since that is the only way you will be able to pass the exams. Copying and other forms of cheating will be dealt with harshly.